# Vamos: Versatile Action Models for Video Understanding

### v2

@@ Abstract
What makes good representations for video understanding, such as anticipating future activities, or answering video-conditioned questions? While earlier approaches focus on end-to-end learning directly from video pixels, we propose to revisit text-based representations, such as general-purpose video captions, which are interpretable and can be directly consumed by large language models (LLMs). Intuitively, different video understanding tasks may require representations that are complementary and at different granularity. To this end, we propose versatile action models (Vamos), a learning framework powered by a large language model as the “reasoner”, and can flexibly leverage visual embedding and free-form text descriptions as its input. To interpret the important text evidence for question answering, we generalize the concept bottleneck model to work with tokens and nonlinear models, which uses hard attention to select a small subset of tokens from the free-form text as inputs to the LLM reasoner. We evaluate Vamos on four complementary video understanding benchmarks, Ego4D, NeXT-QA, IntentQA, and EgoSchema, on its capability to model temporal dynamics, encode visual history, and perform reasoning. Surprisingly, we observe that text-based representations consistently achieve competitive performance on all benchmarks, and that visual embeddings provide marginal or no performance improvement, demonstrating the effectiveness of text-based video representation in the LLM era. We also demonstrate that our token bottleneck model is able to select relevant evidence from free-form text, support test-time intervention, and achieves nearly 5 times inference speedup while keeping a competitive question answering performance. Code and models are publicly released at https://brown-palm.github.io/Vamos/.
@@ 1 Introduction
Figure 1: (a) Visual observations with vastly diverse appearances may be described with the same captions. Our work explores the use of general-purpose text descriptions of video data for action anticipation and question answering. We propose Vamos, a versatile reasoning framework that allows us to study the impact of latent visual representations and free-form text descriptions for downstream applications. (b) Overview of the Vamos framework. It flexibly unifies distributed visual features and textual video representations such as free-form video captions. By performing sequence modeling, Vamos can be applied to multiple video understanding tasks.
@@
We perform extensive evaluations on four benchmarks, including the Ego4D dataset [18] for long-term action anticipation, NeXT-QA [75] and IntentQA [39] for video question answering, and EgoSchema [50] for zero-shot long-form video question answering. We observe that for the direct application of Vamos in the action anticipation task, the text-based representation outperforms its counterpart based on visual embeddings. We further observe that free-form video descriptors serve as an effective long-video representation that generalizes well in zero-shot setting, outperforming the strongest video-language model [71] by 50%. We then confirm that our observations are general, that text-based representation consistently provides competitive performance across all tasks, and that adding visual embeddings surprisingly results in marginal or no performance gains. Finally, we demonstrate that the token bottleneck model is able to select semantically relevant evidence for question answering, and achieves 5x speedup at inference time while maintaining the question answering accuracy.
@@ 2 Related Work
Vision-Language Foundation Models. Vision-Language Models (VLMs) represent a pivotal convergence in artificial intelligence, aiming to bridge the gap between visual understanding and natural language comprehension. Models such as CLIP [59], ALIGN [30], CLOOB [17] bridge the vision and language modalities by learning a text encoder and an image encoder jointly with a contrastive loss, using large datasets consisting of image and caption pairs. Another line of vision-language models [43, 10, 45, 63] use a combination of Masked-Language Modeling (MLM) and Image-Text Matching (ITM) objectives to align specific parts of images with text and enable various downstream tasks such as visual question answering, visual commonsense reasoning, text-based image retrieval, and text-guided object detection. For videos, VIOLET [16] trains an end-to-end transformer for video and language modalities, by representing videos as visual tokens and performing joint masked token modeling. To perform visual-language joint training, speech transcripts are often used as the language modality for videos [64, 16, 86, 85]. The objectives can be combined [82] and the encoders for different modalities can be shared [69]. Compared to existing VLMs, Vamos imposes an “information bottleneck” when text-based representation is used: It converts visual inputs into discrete action labels and free-form text descriptions.
@@ 3.3 Token Bottleneck Models
To implement TBM, we design a lightweight token selector as an add-on module for Vamos (Figure 2 middle). The token selector takes the tokenized embeddings for the text-based video representations $\mathbf{z}_{\text{tvr}}$ and the task-specific tokens $\mathbf{z}_{\text{task}}$ as its inputs. It learns to pick a single token among the candidate tokens $\mathbf{z}_{\text{tvr}}$ by optimizing the objective for a given downstream task. To select a sequence of tokens, we assume the important information is even distributed across the input sequence, and uniformly divide the input sequence into $k$ segments $\{\mathbf{z}_{\text{tvr}}^{(1)}$,…, $\mathbf{z}_{\text{tvr}}^{(k)}\}$, each of which contains $n$ tokens. Each segment $\mathbf{z}_{\text{tvr}}^{(i)}=\{z_{1}^{(i)},...,z_{n}^{(i)}\}$ is fed into the token selector, from which one token $z^{(i)}$ is selected for the task $\mathbf{z}_{\text{task}}$.
@@ 4.1 Task and Datasets
EgoSchema [50] is a recently proposed diagnostic datasets derived from Ego4D for long-form video understanding. Egoschema contains videos that span 3-minute long with high temporal certificate to testify the video understanding model’s capability to understand long videos. It contains 5031 videos and each video has a multiple choice question. It requires zero-shot evaluation and releases answers to 500 questions for development.
@@ 4.2 Implementation
Generating Video Captions. We generate zero-shot, task-agnostic video captions using BLIP-2 for Ego4D LTA, IntentQA, and EgoSchema. We use LLaVA-1.5 to generate captions for NeXT-QA due to its better performance (see Table 4.5 for an ablation study). For Ego4D LTA, we sample the center frame for each video segment to generate its caption. The captions for the 8 observed segments are then concatenated as the video representations. For VQA benchmarks, we first uniformly sample a fixed number of frames for each video, and then generate and concatenate the frame-level captions. We sample 6, 6, and 12 frames for NeXT-QA, IntentQA, and EgoSchema, respectively.
@@ 4.5 Token Bottleneck Models
VQA Performance. We first tokenize and embed the input sequence $\mathbf{x}_{t}$ to obtain $\mathbf{z}_{t}$, which is uniformly partitioned into $k=$ 20 or 40 segments. TBM is then applied on each segment to select one token. In this way, we condense a long sequence (644 tokens on average for NeXT-QA) into $k$ tokens. TBM is jointly optimized with Vamos during training.
For NeXT-QA dataset, we follow previous setting to use captions from 6 frames generated by LLaVA-1.5. For comparison, we also show the vision-only performance taking in 12 frames and the performance with unselected caption-based input. Table 4.5 shows that the TBM leads to an expected performance drop due to discarding over 90% of the input tokens, and increasing $k$ improves the TBM performance. When $k=40$, Vamos achieves a competitive 65.3% accuracy, while only adding 0.7M parameters to the Vamos framework and achieving 5x inference speedup from 1.41s to 0.29s per sample on a single A6000 GPU.
@@
vision all 69.6 67.2 74.7 69.6
text all 75.5 71.3 81.1 75.0
text 20 / 644 59.5 56.8 64.6 59.5
text 40 / 644 65.4 62.4 70.5 65.3
@@
Visualization of the Selected Tokens. TBM not only helps accelerate Vamos’ inference speed, but also allows us to interpret which pieces of evidence are used by the model to make predictions. In Figure 4, we show two positive examples and one negative example of Vamos with TBM from NeXT-QA. The two positive examples demonstrate Vamos’ ability to select tokens that serve as evidence for question answering. The negative example is challenging: even highly related tokens such as “race” and “track” are selected by Vamos, it still fails to reason the purposes of the tyres in the video.
@@ 4.6 Design Choices and Ablation Study
Impact of LLMs on Long-form VideoQA. We study the impact of LLMs on the EgoSchema benchmark. As shown in Table 4.7, GPT-4o achieves significant improvements comparing with GPT-4 and LLaMA2-Chat-7B baselines. This again demonstrates that Vamos can directly benefit from advances in the large language models, making its decoupled perception and reasoning design desirable for practitioners.
@@ 4.7 Comparison with State-of-the-art
Finally, we compare our proposed Vamos with other state-of-the-art models on the four benchmarks in Tables 4.74.74.7, and 4.7. We train Vamos with LLaMA2-7B on Ego4D LTA and LLaMA3-8B on NeXT-QA and IntentQA. On EgoSchema zero-shot VQA, our approach based on GPT-4o outperforms the best vision-language model by 66.8% on the full set. On NeXT-QA, Vamos achieves the best performance and significantly outperforms LLaMA-VQA [36] and SeViLA[83], even though the former uses a LLM with 33B parameters and the latter is trained on additional dataset with temporal localization supervision. On IntentQA, Vamos also outperforms all baselines, with a 28.7% accuracy improvement compared to the best performing prior method. On Ego4D LTA, Vamos outperforms previous works, without relying on domain-specific video encoders [47].
@@
Table 8: Comparison with SOTA on IntentQA
@@ 5 Conclusion
We study different forms of video representations and propose versatile action models (Vamos) as a unified framework to utilize visual- and text-based representations for video understanding. We conduct extensive experiments on long-term action anticipation and video question answer benchmarks. Surprisingly, we observe that direct applications of free-form, general-purpose text-based video representations, such as captions, serve as strong video representation for all benchmarks we consider. Vamos utilizes large language models to perform zero-shot reasoning, and incorporate human feedback via test-time intervention. We further propose the token bottleneck models, which allow the users to interpret the evidence selected by Vamos, and speed up inference by nearly 5x. Vamos achieves state-of-the-art results on Ego4D LTA, IntentQA, NeXT-QA, and outperforms the best vision-language model by over 66% on EgoSchema.
@@
Acknowledgements: This work is supported by Honda Research Institute USA and Samsung Advanced Institute of Technology. We would like to thank Karttikeya Mangalam, Raiymbek Akshulakov, and Shyamal Buch for their kind help with EgoSchema and ATP; Apoorv Khandelwal, Calvin Luo, David Isele, Songpo Li, and Tian Yun for their useful feedback and discussions. Our research was conducted using computational resources at the Center for Computation and Visualization at Brown University.

### v3

@@ Abstract
What makes good representations for video understanding, such as anticipating future activities, or answering video-conditioned questions? While earlier approaches focus on end-to-end learning directly from video pixels, we propose to revisit text-based representations, such as general-purpose video captions, which are interpretable and can be directly consumed by large language models (LLMs). Intuitively, different video understanding tasks may require representations that are complementary and at different granularity. To this end, we propose versatile action models (Vamos), a learning framework powered by a large language model as the “reasoner”, and can flexibly leverage visual embedding and free-form text descriptions as its input. To interpret the important text evidence for question answering, we generalize the concept bottleneck model to work with tokens and nonlinear models, which uses hard attention to select a small subset of tokens from the free-form text as inputs to the LLM reasoner. We evaluate Vamos on five complementary benchmarks, Ego4D, NeXT-QA, IntentQA, Spacewalk-18, and EgoSchema, on its capability to model temporal dynamics, encode visual history, and perform reasoning. Surprisingly, we observe that text-based representations consistently achieve competitive performance on all benchmarks, and that visual embeddings provide marginal or no performance improvement, demonstrating the effectiveness of text-based video representation in the LLM era. We also demonstrate that our token bottleneck model is able to select relevant evidence from free-form text, support test-time intervention, and achieves nearly 5 times inference speedup while keeping a competitive question answering performance. Code and models are publicly released at https://brown-palm.github.io/Vamos/.
@@ 1 Introduction
Figure 1: (a) Visual observations with vastly diverse appearances may be described with the same captions. Our work explores the use of general-purpose text descriptions of video data for action anticipation and question answering. We propose Vamos, a versatile reasoning framework that allows us to study the impact of latent visual representations and free-form text descriptions for downstream applications. (b) Overview of the Vamos framework. It flexibly unifies distributed visual features and text-based representations such as video captions, and can be applied to diverse video understanding tasks.
@@
We perform extensive evaluations on five benchmarks, including the Ego4D dataset [17] for long-term action anticipation, NeXT-QA [75] and IntentQA [39] for video question answering, Spacewalk-18 [37] for long-form procedural video understanding, and EgoSchema [50] for zero-shot long-form video question answering. We observe that for the direct application of Vamos in the action anticipation task, the text-based representation outperforms its counterpart based on visual embeddings. We further observe that free-form video descriptors serve as an effective long-video representation that generalizes well in zero-shot setting, outperforming the strongest video-language model [71] by 66%. We then confirm that our observations are general, that text-based representation consistently provides competitive performance across all tasks, and that adding visual embeddings surprisingly results in marginal or no performance gains. Finally, we demonstrate that the token bottleneck model is able to select semantically relevant evidence for question answering, and achieves 5x speedup at inference time while maintaining the question answering accuracy.
@@ 2 Related Work
Vision-Language Foundation Models. Models such as CLIP [59] and ALIGN [29] bridge the vision and language modalities by learning a text encoder and an image encoder jointly with a contrastive loss on image and caption pairs. Another line of vision-language models [43, 10, 45, 63] combines the masked language modeling objective with image-text contrastive learning, and focus on downstream tasks such as visual question answering, visual commonsense reasoning, and text-guided object detection. For videos, VIOLET [16] trains an end-to-end transformer for video and language modalities, by representing videos as visual tokens and performing joint masked token modeling. To perform visual-language joint training, speech transcripts are often used as the language modality for videos [64, 16, 86, 85]. The objectives can be combined [82] and the encoders for different modalities can be shared [69]. Compared to existing VLMs, Vamos imposes an “information bottleneck” when text-based representation is used: It converts visual inputs into discrete action labels and free-form text descriptions.
@@ 3.3 Token Bottleneck Models
To implement TBM, we design a lightweight token selector as an add-on module for Vamos (Figure 2 middle). It takes the tokenized embeddings for the text-based video representations $\mathbf{z}_{\text{tvr}}$ and the task-specific tokens $\mathbf{z}_{\text{task}}$ as its inputs. It learns to pick a single token among the candidate tokens $\mathbf{z}_{\text{tvr}}$ by optimizing the objective for a given downstream task. To select a sequence of tokens, we assume the important information is even distributed across the input sequence, and uniformly divide the input sequence into $k$ segments $\{\mathbf{z}_{\text{tvr}}^{(1)}$,…, $\mathbf{z}_{\text{tvr}}^{(k)}\}$, each of which contains $n$ tokens. Each segment $\mathbf{z}_{\text{tvr}}^{(i)}=\{z_{1}^{(i)},...,z_{n}^{(i)}\}$ is fed into the token selector, from which one token $z^{(i)}$ is selected for the task $\mathbf{z}_{\text{task}}$.
@@ 4.1 Task and Datasets
EgoSchema [50] is annotated on Ego4D videos for long-form video QA. Each video is around 3 minutes and the temporal certificate for humans to solve each task is around 100 seconds. It has 5,031 videos and each video is annotated with a multiple-choice question. All examples are for zero-shot evaluation.
Spacewalk-18 [37] is a long-form procedural video understanding benchmark collected on 18 spacewalk videos. The total duration is over 96 hours. We evaluate on the step recognition task which has a temporal certificate of 140 seconds. We follow the zero-shot setup with 1-minute context window, and report step recognition accuracy on the test set.
@@ 4.2 Implementation
Generating Video Captions. We generate zero-shot, task-agnostic video captions using BLIP-2 for Ego4D LTA, IntentQA, and EgoSchema. We use LLaVA-1.5 to generate captions for NeXT-QA and Spacewalk-18 due to its better performance (see Table 4.5). For Ego4D LTA, we sample the center frame for each video segment to generate its caption. The captions for the 8 observed segments are then concatenated as the video representations. For VQA benchmarks, we first uniformly sample a fixed number of frames for each video, and then generate and concatenate the frame-level captions. We sample 6, 6, 12, and 12 frames for NeXT-QA, IntentQA, Spacewalk-18, and EgoSchema, respectively. For Spacewalk-18, we additional use its provided speech narrations.
@@ 4.5 Token Bottleneck Models
VQA Performance. We first tokenize and embed the input sequence $\mathbf{x}_{t}$ to obtain $\mathbf{z}_{t}$, which is uniformly partitioned into $k=$ 20 or 40 segments. TBM is then applied on each segment to select one token. In this way, we condense a long sequence (644 tokens on average for NeXT-QA) into $k$ tokens. TBM is jointly optimized with Vamos with LLaMA-3 during training.
For NeXT-QA dataset, we follow previous setting to use captions from 6 frames generated by LLaVA-1.5. For comparison, we also show the vision-only performance taking in 12 frames and the performance with unselected caption-based input. Table 4.5 shows that the TBM leads to an expected performance drop due to discarding over 90% of the input tokens, and increasing $k$ improves the TBM performance. When $k=40$, Vamos achieves a competitive 69.6% accuracy, while only adding 0.7M parameters to the Vamos framework and achieving 5x inference speedup from 1.41s to 0.29s per sample on a single A6000 GPU.
@@
vision all 71.9 67.4 75.6 71.0
text all 77.2 75.3 81.7 77.3
text 20 / 644 68.1 64.6 70.8 67.4
text 40 / 644 70.1 67.4 72.2 69.6
@@
Visualization of the Selected Tokens. TBM not only helps accelerate Vamos’ inference speed, but also allows us to interpret which pieces of evidence are used by the model to make predictions. In Figure 4, we show two positive examples and one negative example of Vamos with TBM from NeXT-QA. The two positive examples demonstrate Vamos’ ability to select tokens that serve as evidence for question answering. The negative example is challenging: even highly related tokens such as “snow” and “glide” are selected by Vamos, it still fails to reason the correct activities in the video.
@@ 4.6 Design Choices and Ablation Study
Impact of LLMs on Long-form VideoQA. We study the impact of LLMs on the EgoSchema benchmark. As shown in Table 4.7, GPT-4o achieves significant improvements comparing with GPT-4 and LLaMA2-Chat-7B baselines, which again demonstrates that Vamos can directly benefit from advances in LLMs.
@@ 4.7 Comparison with State-of-the-art
We compare our proposed Vamos with other state-of-the-art models in Tables 4.74.74.7, and 4.7. We train Vamos with LLaMA2-7B on Ego4D LTA and LLaMA3-8B on NeXT-QA and IntentQA. On EgoSchema zero-shot VQA, our approach based on GPT-4o outperforms the best vision-language model by 66.8% on the full set. On NeXT-QA, Vamos achieves the best performance and significantly outperforms LLaMA-VQA [35] and SeViLA[83], even though the former uses a LLM with 33B parameters and the latter is trained on additional dataset with temporal localization supervision. On IntentQA, Vamos also outperforms all baselines, with a 28.7% accuracy improvement compared to the best performing prior method. On Ego4D LTA, Vamos outperforms previous works, without relying on domain-specific video encoders [47]. Finally, on Spacewalk-18, we use GPT-4o as the LLM and achieves 18.6% accuracy, which significantly outperforms the prior best zero-shot performance of 13.6%.
@@
Table 8: Comparison with SOTA on IntentQA.
@@ 5 Conclusion
We study different forms of video representations and propose versatile action models (Vamos) as a unified framework to utilize visual- and text-based representations for video understanding. We conduct extensive experiments on long-term action anticipation and video question answer benchmarks. Surprisingly, we observe that direct applications of free-form, general-purpose text-based video representations, such as captions, serve as strong video representation for all benchmarks we consider. Vamos utilizes large language models to perform zero-shot reasoning, and incorporate human feedback via test-time intervention. We further propose the token bottleneck models, which allow the users to interpret the evidence selected by Vamos, and speed up inference by nearly 5x. Vamos achieves state-of-the-art results on Ego4D LTA, IntentQA, NeXT-QA, Spacewalk-18, and outperforms the best vision-language model by over 66% on EgoSchema.
@@
Acknowledgements
This work is supported by Honda Research Institute USA and Samsung Advanced Institute of Technology. We would like to thank Karttikeya Mangalam, Raiymbek Akshulakov, and Shyamal Buch for their kind help with EgoSchema and ATP; Apoorv Khandelwal, Calvin Luo, David Isele, Songpo Li, and Tian Yun for their useful feedback and discussions. Our research was conducted using computational resources at the Center for Computation and Visualization at Brown University.
Appendix 0.A Additional Experiments and Results
Textual representation format. On the Ego4D LTA task, we observe that explicitly regularizing the text-based inputs to contain only verbs and nouns lead to slightly improved performance (0.878 versus 0.890 action edit distance). We hypothesize that this is due to the nature of the task, which focuses solely on predicting future verbs and nouns. In Table 0.A, we again compare the action-based and caption-based text representations, but on the EgoSchema zero-shot VQA task. In contrast to the LTA performance, the caption-based representation performs much better, intuitively because solving the video question answering would require more details about the video, which can be provided by the general-purpose captions.
Frame number on EgoSchema. In Table 3 we compare different frame numbers and the zero-shot performance on EgoSchema. We uniformly sample 1, 4, and 12 frames from the videos and concatenate the captions of each frame to form the text video representations. We use GPT-3.5 as the reasoner. The results show that using more frames leads to better performance as more information and temporal evidence are provided.
Temporal ordering information. We investigate the influence of temporal information for long-form video understanding by shuffling the frame order when concatenating captions. In Table 0.A, we compare the ordered and shuffled 12-frame captions. Surprisingly, the performance drop (2%) by shuffling the video captions is not as significant as we expect. We suspect that LLMs may have strong capability of “auto-correcting” the order of the input sentences, even if they are shuffled.
Table A1: Comparison of action and caption for zero-shot VQA on Egoschema.
Model Input Type Full Set Acc.
GPT-4 action 38.12%
GPT-4 caption 48.26%
Table A2: Ablation on the influence of temporal information. 12 frames are sampled.
Shuffle   Full Set Acc.
39.22%
41.24%
Appendix 0.B Addition Implementation Details
Vamos. We train Vamos on 4 A6000 GPUs for 2, 5, 10 epochs on Ego4D, IntentQA, and NeXT-QA respectively. For NeXT-QA, we set the maximum sequence length of pure vision input as 128, and 1200 for captions and captions + vision inputs, for IntentQA, the sequence lengths are set as 128 and 512 respectively.
Token bottleneck model. For the token bottleneck model (TBM), we use a 2-layer transformer encoder with 2 attention heads, and a hidden size of 128. No additional positional encoding is added by the token bottleneck model. In order to condense the input sequence for interpretability and select the most relevant tokens, the TBM is task dependent, namely taking the questions and the candidate answers (when available) as inputs along with the video captions. The selected tokens, as opposed to their corresponding encoded embeddings by the token bottleneck, are fed into the LLM, as illustrated in Figure 2 (middle) in the main paper.
Appendix 0.C Additional Visualization
Figure A3 illustrates the input format for long-term action anticipation and video question answering tasks. In the LTA task, the task-specific inputs are the discrete action labels and the target output is a future action sequence. For supervised VQA, the task-specific input is composed of instructions, video representations (vision and text), question and choices. The target output is the chosen answer to the question. For zero-shot VQA on EgoSchema, the video representation only consists of the text descriptions. Figure A3 shows the prompt designs for all tasks.
Figure A4 shows four example video captions generated by LLaVA and BLIP-2 on the NeXT-QA dataset, respectively. We observe that in general the LLaVA-generated captions are longer and more detailed. Recall that Table 5 shows that the more detailed LLaVA captions also lead to higher VQA performance.
In Figure A1 and Figure A2, we provide additional prediction examples from Vamos and Vamos with the token bottleneck model respectively.
Figure A1: More examples of Vamos video question answering and manual intervention.
Figure A2: Additional illustration of predictions with TBM and manual intervention. The token bottleneck selects tokens highly related to the question and intervention corrects wrong predictions without training.
Figure A3: Input of Vamos for long-term action anticipation and video question answering. <v> in prompts denotes vision tokens.
Figure A4: Example of video captions generated by BLIP-2 and LLaVA on the NeXT-QA dataset, respectively.